The length of a rectangular piece of land is 90 yards more than three times its width. The perimeter is 780 yards. find its dimensions.
Answer by Cromlix(4381)

Width = x
Length = 90 + 3x
Perimeter = 2*length + 2*width
780 = 2(90 + 3x) + 2x
780 = 180 + 6x + 2x
780 - 180 = 8x
8x = 600
x = 75yds
Length = 90 + 3(75)
= 315yds
Width = 75yds
Hope this helps.
:-)